# Fractional Programming for Stochastic Precoding over Generalized Fading Channels

### v1

@@ Abstract
Index Terms:
Stochastic precoding, multiple-input multiple-output (MIMO), first and second moments, long-term average weighted sum rates, matrix fractional programming (FP).
@@ I Introduction
This paper pursues the maximization of the long-term average weighted sum rates. For this problem, [26] proposes a data-driven stochastic precoding algorithm that takes a large number of fading channel samples as training data. Each sample corresponds to a deterministic precoding subproblem—which is solvable by the weighted minimum mean square error (WMMSE) algorithm [33, 34]; the main idea of [26] is to iteratively average out the deterministic precoding solutions. As shown in [26], the iterations must converge to a stationary point of the original problem provided that the number of fading channel samples is sufficiently large. This online optimization method has evolved in some more recent works, e.g., [29] addresses the deterministic subproblems via neural networks, [27] combines the deterministic solutions in a proximal fashion, and [13] tailors the online optimization to the Gaussian case. In contrast, many other works propose the offline optimization methods. However, even when fading channels are assumed to be Gaussian, the expectation of data rate is still difficult to tackle directly, so [12] proposes a heuristic approximation of the original objective function, while [8, 32] suggest using an upper bound to approximate it. Nevertheless, since this is a maximization problem, it is more desirable to find a lower bound for approximation. Toward this end, [11] constructs a lower bound on the expectation of weighted sum rates, but the bound is difficult to compute in practice. Some computable lower bounds are proposed in [9, 10] for the single-cell network with a single data stream to each user. We remark that the above bounds, either upper or lower, are all designed for the Gaussian fading channel model. In contrast, the lower bound proposed in this paper accounts for a generic fading channel model which need not be Gaussian.
@@ II System Model and Problem Formulation
Consider an $L$-cell downlink MIMO network, where each cell consists of one BS and $K$ user terminals. We refer to the BS of the $\ell$th cell as BS $\ell$, and the $k$th user terminal in the $j$th cell as user $(j,k)$, for $\ell,j=1,2,\ldots,L$ and $k=1,2,\ldots,K$. Assume that each BS has $M_{t}$ transmit antennas while each user terminal has $M_{r}$ receive antennas. Let $\mathbf{H}_{jk,\ell}\in\mathbb{C}^{M_{r}\times M_{t}}$ be the block fading channel from BS $\ell$ to user $(j,k)$. These $\mathbf{H}_{jk,\ell}$’s are modeled as i.i.d. random variables across the different blocks. We assume that the distribution function of $\mathbf{H}_{jk,\ell}$ is unknown or cannot be written analytically; only the first and second moments of random channels are available:
@@
The goal of stochastic precoding is to design a set of precoding matrices $\{\mathbf{V}_{jk}\in\mathbb{C}^{M_{t}\times M_{r}}:\forall(j,k)\}$ that is “good” for all the blocks on average.
@@
with $\sigma^{2}$ denoting the background noise power. We aim to maximize the weighted sum rates across many blocks; by the law of large numbers, the problem can be written as
@@ III-A Stochastic Extension of Matrix FP
We start by showing how the quadratic transform and the Lagrangian dual transform in [36] are extended for the stochastic case. To avoid complication in notation, and to present the results in full generality, we consider the following generic problem formulation. Let $\mathbf{A}_{n}(x\mid\theta^{+}_{n})\in\mathbb{S}^{M\times M}_{+}$ and $\mathbf{B}_{n}(x\mid\theta^{-}_{n})\in\mathbb{S}^{M\times M}_{++}$ be a pair of matrix-valued functions of $x$, for $n=1,2,\ldots,N$, where $\{\theta^{+}_{n},\theta^{-}_{n}\}$ is a set of implicit random variables that impact the function outputs. To ease notation, the arguments $(x\mid\theta^{+}_{n})$ and $(x\mid\theta^{-}_{n})$ are omitted in the rest of this section, unless we wish to emphasize their specific values. Let us consider an abstraction of the stochastic precoding problem:
@@
where $\mathcal{X}$ is a nonempty constraint set and each $\omega_{n}\geq 0$ is a positive weight. Following the Lagrangian dual transform [36], we try to move the matrix ratio term out of the logdet function, as stated in the following theorem.
@@ III-B Proposed Lower Bound
subject to $\quad\bm{\Gamma}_{n}\in\mathbb{S}^{M\times M}_{+},\;\forall n$ (16b)
@@
$\hat{f}(x,\bm{\Gamma},\mathbf{Y})$ $\leq\sup_{\bm{\Gamma},\mathbf{Y}}\hat{f}(x,\bm{\Gamma},\mathbf{Y})$ (19)
$=\sup_{\bm{\Gamma},\mathbf{Y}}\mathbb{E}[g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+}_{n},\theta^{-}_{n})]$ (20)
$\leq\mathbb{E}\bigg{[}\sup_{\bm{\Gamma},\mathbf{Y}}g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+}_{n},\theta^{-}_{n})\bigg{]}$ (21)
$=f(x).$ (22)
@@ IV FP-Based Stochastic Precoding
$\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})=\\ \sum_{(j,k)}\omega_{jk}\Bigg{[}\log\big{|}\mathbf{I}+\mathbf{\Gamma}_{jk}\big{|}-\mathrm{Tr}\big{(}\mathbf{\Gamma}_{jk}\big{)}+\mathrm{Tr}\Bigg{(}\big{(}\mathbf{I}+\mathbf{\Gamma}_{jk}\big{)}\Bigg{(}2\Re\big{\{}\mathbf{V}_{jk}^{H}\mathbf{C}_{jk}^{H}\mathbf{Y}_{jk}\big{\}}-\mathbf{Y}_{jk}^{H}\Bigg{(}\sum_{(\ell,s)}\mathbf{U}_{jk,\ell s}+\sigma^{2}\mathbf{I}\Bigg{)}\mathbf{Y}_{jk}\Bigg{)}\Bigg{)}\Bigg{]}$ (23)
@@
where the new objective function $\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})$ is shown in (23) at the top of next page, with the shorthand
@@
×$\qquad\times(\mathbf{D}_{jk,\ell})_{(m^{\prime}-1)M_{r}+m,(n^{\prime}-1)M_{r}+n}\Big{)}$ (28)
@@
×$\qquad\qquad\qquad\quad\times\mathbb{E}\big{[}\big{(}\mathbf{H}^{c}_{\ell s,j}\big{)}_{m^{\prime},m}\big{(}\mathbf{H}_{\ell s,j}\big{)}_{n^{\prime},n}\big{]}\Big{)}$
@@
×$\qquad\qquad\qquad\times(\mathbf{D}_{\ell s,j}^{c})_{(m-1)M_{r}+m^{\prime},(n-1)M_{r}+n^{\prime}}\Big{)}.$ (34)
@@ V Acceleration for Large-Scale MIMO
$\mathbf{V}^{\star}_{jk}=\left\{\begin{array}[]{ll}{\bm{\Psi}}_{jk}&\text{if}\;\sum^{K}_{k=1}\mathrm{Tr}(\bm{\Psi}_{jk}^{H}\bm{\Psi}_{jk})\leq P\vspace{0.5em}\\ \beta_{j}\bm{\Psi}_{jk}&\text{otherwise},\end{array}\right.$ (51)
@@
We only sketch the basic ideas which follow the existing work [40] closely. First, Algorithm 2 can be recognized as a minorization-maximization (MM) method for problem (47); because each step of Algorithm 2 boils down to solving a convex subproblem, the convergence to a stationary point can be verified by the MM theory [40]. Moreover, using the inequality in Lemma 1 to bound $\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})$ can be thought of as constructing a surrogate function for problem (26), so each stationary point of problem (47) can be mapped to a stationary point of problem (26) according to the MM theory. ∎
@@ VI Simulations
Figure 1: Lower bounds v.s. original objective for different noise power levels.
Figure 2: Lower bounds v.s. original objective for different values of the temporal correlation coefficient $\rho_{jk}$.
@@ VI-A Settings
where $\bar{\mathbf{H}}_{jk,\ell}$ is the static component, $\mathbf{W}_{jk,\ell}\odot\mathbf{X}_{jk,\ell}$ is the random component, and $\rho_{jk}\in(0,1]$ is the tradeoff factor. These parameters are further specified in the following. The real matrix $\mathbf{W}_{jk,\ell}\in\mathbb{R}^{M_{r}\times M_{t}}$ reflects on the large-scale fading, each entry of which is computed as $10^{-\xi/10}$ where $\xi=0.5\times(128.1+37.6\log_{10}d+\tau)$, $d$ denoting the distance in km between the corresponding pair of BS and user, $\tau$ being a zero-mean Gaussian random variable with a standard deviation of 8 dB for the shadowing effect. The fixed matrix $\bar{\mathbf{H}}_{jk,\ell}$ is a realization of $\mathbf{W}_{jk,\ell}\odot\mathbf{X}_{jk,\ell}$ at the beginning of the test. The random complex matrix $\mathbf{X}_{jk,\ell}\in\mathbb{C}^{M_{r}\times M_{t}}$ has each entry independently drawn from the standard complex Gaussian distribution $\mathcal{CN}(0,1)$. The tradeoff factor $\rho_{jk}$, a.k.a. the temporal correlation coefficient, is generated by Jake’s model [41]. We consider $1000$ independent blocks and evaluate their average rates.
@@
Figure 3: The long-term average sum rates achieved by the different algorithms under the different fading models and under the different noise power levels.
We further specify the network setting. The radius of each cell equals $300$ meters. We let $P=30$ dBm and $\omega_{jk}=1$. The noise power is $-90$ dBm by default. For the multi-cell case, we consider 7 wrapped-around cells. We let $M_{t}=64$, $K=32$, and $M_{r}=2$ by default for the single-cell case, and $M_{t}=32$, $K=16$, and $M_{r}=2$ by default for the multi-cell case. Moreover, the benchmark methods for stochastic precoding include:
VI-B Test Results
We begin by evaluating the tightness of the proposed lower bound in (3). Fig. 1 and Fig. 2 consider the Rayleigh fading model. We remark that the lower bound in [11] is not included in the above two figures because it is uncomputable. It can be seen that the proposed bound is much tighter than the existing one in [12].
Next, we compare the expected sum rates achieved by different algorithms under the different values of noise power $\sigma^{2}$ in Fig. 3. Note that RLPD is limited to the single-cell Rayleigh fading case, so it only appears in Fig. 3(a). For SWMMSE, we provide a total of $1000$ channel samples for online optimization. Because Algorithm 1 and Algorithm 2 only differ in the computational complexity, they produce the same performance curves in Fig. 3. Observe that the WMMSE algorithm that only depends on the static components of fading channels lead to the worst performance. The proposed method attains much better performance, which is about 30% higher than that of WMMSE. Observe also that SWMMSE yields fairly nice performance in the single-cell case, which is almost equally good as the proposed method. However, SWMMSE becomes much worse in the presence of multiple cells since then it requires many more channel samples. In practice, online optimization by SWMMSE can be quite costly. RLPD has good performance, but unfortunately it is limited to the single-cell Rayleigh fading case.
@@ VI-B Test Results
Figure 4: The convergence of the different algorithms in terms of the number of iterations under the different fading models.
(a) Single-cell Rayleigh fading case
(b) Multi-cell Rayleigh fading case
(c) Single-cell Nakagami-$m$ fading case
(d) Multi-cell Nakagami-$m$ fading case
@@
Figure 5: The convergence of the different algorithms in terms of elapsed time under the different fading models.

### v2

@@ I Introduction
This paper pursues the maximization of the long-term average weighted sum rates. Here and throughout, “long-term” refers to the timescale in which the first-order and second-order moments are evaluated for the fading channels. In other words, the data rate can be approximated as an ergodic process under this long-term setting. For this problem, [23] proposes a data-driven stochastic precoding algorithm that takes a large number of fading channel samples as training data. Each sample corresponds to a deterministic precoding subproblem—which is solvable by the weighted minimum mean-squared error (WMMSE) algorithm [5, 32]; the main idea of [23] is to iteratively average out the deterministic precoding solutions. As shown in [23], the iterations must converge to a stationary point of the original problem provided that the number of fading channel samples is sufficiently large. This online optimization method has evolved in some more recent works, e.g., [14] addresses the deterministic subproblems via neural networks, [43] combines the deterministic solutions in a proximal fashion, and [39] tailors the online optimization to the Gaussian case. In contrast, many other works propose the offline optimization methods. However, even when fading channels are assumed to be Gaussian, the expectation of data rate is still difficult to tackle directly, so [4] proposes a heuristic approximation of the original objective function, while [36, 34] suggest using an upper bound to approximate it. Nevertheless, since this is a maximization problem, it is more desirable to find a lower bound for approximation. Toward this end, [16] constructs a lower bound on the expectation of weighted sum rates, but the bound is difficult to compute in practice. Some computable lower bounds are proposed in [31, 30] for the single-cell network with a single data stream to each user. We remark that the above bounds, either upper or lower, are all designed for the Gaussian fading channel model. In contrast, the lower bound proposed in this paper accounts for a generic fading channel model which need not be Gaussian.
@@ II System Model and Problem Formulation
Consider an $L$-cell downlink MIMO network, where each cell consists of one BS and $K$ user terminals. We refer to the BS of the $j$th cell as BS $j$, and the $k$th user terminal in the $j$th cell as user $(j,k)$, for $j,\ell=1,2,\ldots,L$ and $k=1,2,\ldots,K$, where $j$ is the main index for BS while $\ell$ is an ancillary index. Assume that each BS has $M_{t}$ transmit antennas while each user terminal has $M_{r}$ receive antennas. Let $\mathbf{H}_{jk,\ell}\in\mathbb{C}^{M_{r}\times M_{t}}$ be the block fading channel from BS $\ell$ to user $(j,k)$ in $j$th cell. These $\mathbf{H}_{jk,\ell}$’s are modeled as i.i.d. random variables across the different blocks. We assume that the distribution function of $\mathbf{H}_{jk,\ell}$ is unknown or cannot be written analytically; only the first and second moments of random channels are available:
@@
How these moment statistics can be obtained has been extensively discussed in the literature [42, 24, 48]. Moreover, as pointed out in the more recent work [18], the channels can be predicted by the modern tools of artificial intelligence (AI) and digital twins; in particular, [18] shows that the channel statistics (e.g., the first-order and second-order moments) are much easier to predict than the specific channel realization in real-world networks. The goal of stochastic precoding is to design a set of precoding matrices $\{\mathbf{V}_{jk}\in\mathbb{C}^{M_{t}\times M_{r}}:\forall(j,k)\}$ that is “good” for all the blocks on average.
@@
with $\sigma^{2}$ denoting the background noise power. We use two distinct BS indices $j$ and $\ell$ to denote the cross-cell channels, e.g., the interference channel from the neighboring BS $\ell$ to the $k$th user in cell $j$ in (4) is denoted by $\mathbf{H}_{jk,\ell}$. We aim to maximize the long-term weighted sum rates across many blocks. By the law of large numbers, the problem can be written as
@@ III-A Stochastic Extension of Matrix FP
[…]
@@
where $\mathcal{X}$ is a nonempty constraint set and each $\omega_{n}\geq 0$ is a positive weight. Following the Lagrangian dual transform [25], we try to move the matrix ratio term out of the log-det function, as stated in the following theorem.
@@
However, Theorem 1 and Theorem 2 alone are insufficient for solving the stochastic beamforming problem, because their auxiliary variables are difficult to decide. To address this issue, we further propose a lower bound in Section III-B, which is a key result of this paper.
@@ III-B Proposed Lower Bound
subject to $\quad\bm{\Gamma}_{n}\in\mathbb{S}^{M\times M}_{+},\;\forall n,$ (16b)
@@
We first show that the interchange of $\mathbb{E}$ and $\sup$ in (17) is mathematically well-posed. Since the feasible set of $(\bm{\Gamma},\mathbf{Y})$ needs to be nonempty, and that the objective function $g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+},\theta^{-})$ is upper bounded, so $\sup_{\mathbf{\Gamma},\mathbf{Y}}g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+},\theta^{-})$ must exist. Note also that $g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+},\theta^{-})$ is a continuous function defined on the measurable domain $(x,\mathbf{\Gamma},\mathbf{Y})\in\mathcal{X}\times\mathbb{S}_{+}^{M\times M}\times\mathbb{C}^{M\times M}$, so $g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+},\theta^{-})$ is a measurable function, and its supremum is a measurable function as well. As a result, $g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+},\theta^{-})$ and $\sup_{\mathbf{\Gamma},\mathbf{Y}}g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+},\theta^{-})$ are both Lebesgue integrable, so $\mathbb{E}[\sup_{\mathbf{\Gamma},\mathbf{Y}}g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+},\theta^{-})]$ and $\mathbb{E}[g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+},\theta^{-})]$ must exist. Further, since $\mathcal{X}$ is nonempty, $\sup_{\mathbf{\Gamma},\mathbf{Y}}\mathbb{E}[g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+},\theta^{-})]$ must exist.
@@
$\hat{f}(x,\bm{\Gamma},\mathbf{Y})$ $\leq\sup_{\bm{\Gamma},\mathbf{Y}}\hat{f}(x,\bm{\Gamma},\mathbf{Y})$ (19a)
$=\sup_{\bm{\Gamma},\mathbf{Y}}\mathbb{E}[g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+}_{n},\theta^{-}_{n})]$ (19b)
$\leq\mathbb{E}\bigg[\sup_{\bm{\Gamma},\mathbf{Y}}g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+}_{n},\theta^{-}_{n})\bigg]$ (19c)
$=f(x).$ (19d)
@@
$\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})=\sum_{(j,k)}\omega_{jk}\Bigg[\log\big|\mathbf{I}+\mathbf{\Gamma}_{jk}\big|-\mathrm{Tr}\big(\mathbf{\Gamma}_{jk}\big)+\mathrm{Tr}\Bigg(\big(\mathbf{I}+\mathbf{\Gamma}_{jk}\big)\Bigg(2\Re\big\{\mathbf{V}_{jk}^{H}\mathbf{C}_{jk}^{H}\mathbf{Y}_{jk}\big\}-\mathbf{Y}_{jk}^{H}\Bigg(\sum_{(\ell,s)}\mathbf{U}_{jk,\ell s}+\sigma^{2}\mathbf{I}\Bigg)\mathbf{Y}_{jk}\Bigg)\Bigg)\Bigg]$ (20)
In problem (16), to make the bound $\hat{f}(x,\bm{\Gamma},\mathbf{Y})$ as tight as possible, we maximize $\hat{f}(x,\bm{\Gamma},\mathbf{Y})$ over $\mathbf{V}$ and $\bm{\Gamma}$ to get:
$\tilde{\bm{\Gamma}}^{\star}_{n}$ $=\mathbb{E}[\sqrt{\mathbf{A}_{n}}]^{H}\Big(\mathbb{E}[\mathbf{A}_{n}]+\mathbb{E}[\mathbf{B}_{n}]$
$\qquad-\mathbb{E}[\sqrt{\mathbf{A}_{n}}]\mathbb{E}[\sqrt{\mathbf{A}_{n}}]^{H}\Big)^{-1}\mathbb{E}[\sqrt{\mathbf{A}_{n}}],$ (21)
$\tilde{\mathbf{Y}}_{n}^{\star}$ $=\Big(\mathbb{E}[\mathbf{A}_{n}]+\mathbb{E}[\mathbf{B}_{n}]\Big)^{-1}\mathbb{E}[\sqrt{\mathbf{A}_{n}}],$ (22)
where the arguments of $\mathbf{A}_{n}$ and $\mathbf{B}_{n}$ are omitted to ease notation. Importantly, the above two solutions can be treated as the functions of $x$, so substituting them into $\hat{f}(x,\bm{\Gamma},\mathbf{Y})$ leads to a lower bound $\tilde{f}(x)$ without $\bm{\Gamma}$ and $\mathbf{Y}$, i.e.,
[…]
where the equality holds when $\mathrm{Pr}\{(\theta^{+}_{n},\theta^{-}_{n}):\mathbf{\Gamma}^{\star}_{n}=\tilde{\bm{\Gamma}}^{\star}_{n}\;\text{and}\;\mathbf{Y}_{n}^{\star}=\tilde{\mathbf{Y}}^{\star}_{n}\}=1$, i.e., when $(\mathbf{\Gamma},\mathbf{Y})$ are deterministic functions of moments. As compared to the primal objective $f(x)$, the new bound $\tilde{f}(x)$ looks simpler in that it gets rid of the expectation. However, $\tilde{f}(x)$ is still difficult to optimize directly because of its nonconvexity. If we consider the parameterized bound $\hat{f}(x,\bm{\Gamma},\mathbf{Y})$ instead, then the solution of $x$ can be readily obtained in closed form with the auxiliary variables $(\bm{\Gamma},\mathbf{Y})$ updated iteratively; this can be recognized as a minorization-maximization (MM) procedure with provable performance.
@@ IV FP-Based Stochastic Precoding
where the new objective function $\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})$ is shown in (20), with the shorthand
@@
$\qquad\times(\mathbf{D}_{jk,\ell})_{(m^{\prime}-1)M_{r}+m,(n^{\prime}-1)M_{r}+n}\Big)$ (28)
@@
$\qquad\qquad\qquad\quad\times\mathbb{E}\big[\big(\mathbf{H}^{c}_{\ell s,j}\big)_{m^{\prime},m}\big(\mathbf{H}_{\ell s,j}\big)_{n^{\prime},n}\big]\Big)$
@@
$\qquad\qquad\qquad\times(\mathbf{D}_{\ell s,j}^{c})_{(m-1)M_{r}+m^{\prime},(n-1)M_{r}+n^{\prime}}\Big).$ (34)
@@
The following remarks in order point out some subtle aspects of Algorithm 1 and Proposition 1.
Remark 1
The proposed algorithm guarantees a stationary point of the lower bound but not the original problem. We also wish to point out that the above result is by no means trivial. In the existing literature, [16] show that their algorithms converge to stationary points of the new problems, but the connection between the new problems and the original problem cannot be proved. Moreover, [36, 31, 34] use the upper bound to approximate the problem and show that their solutions are the stationary points of the new problem. However, since the original problem is a maximization problem, it makes more sense to consider a lower bound as in our work. To the best of our knowledge, [23] and [43] are the only works that guarantee stationary points of the original problem, but they are in essence based on Monte Carlo method and require a huge set of instantaneous CSI for the training purpose.
Remark 2
The proposed method is in essence to approximate the fading channels as Gaussian. After all, with only the first-order and second-order moments available, it is most reasonable to adopt the Gaussian approximation. Nevertheless, it is not that trivial to show that the Gaussian approximation can yield a lower bound for this particular problem case.
@@ V Acceleration for Large-Scale MIMO
Remark 4
The gap between $\hat{f}$ and $\zeta$ can be computed as
$\hat{f}-\zeta=$ $\sum_{(j,k)}\mathrm{Tr}\Big((\mathbf{V}_{jk}-\mathbf{Z}_{jk})^{H}$
$\qquad\times\left(\lambda_{\max}(\mathbf{\Xi}_{j})\mathbf{I}-\mathbf{\Xi}_{j}\right)(\mathbf{V}_{jk}-\mathbf{Z}_{jk})\Big)$ (45a)
$\leq$ $\sum_{(j,k)}\lambda_{\max}\left(\lambda_{\max}(\mathbf{\Xi}_{j})\mathbf{I}-\mathbf{\Xi}_{j}\right)$
$\qquad\times\mathrm{Tr}\Big((\mathbf{V}_{jk}-\mathbf{Z}_{jk})^{H}(\mathbf{V}_{jk}-\mathbf{Z}_{jk})\Big)$ (45b)
$=$ $\sum_{(j,k)}\Big(\lambda_{\max}(\mathbf{\Xi}_{j})-\lambda_{\min}(\mathbf{\Xi}_{j})\Big)$
$\qquad\times\mathrm{Tr}\Big((\mathbf{V}_{jk}-\mathbf{Z}_{jk})^{H}(\mathbf{V}_{jk}-\mathbf{Z}_{jk})\Big).$ (45c)
Observe that the gap increases with the spectral spread of $\Xi_{j}$.
@@
$\mathbf{V}^{\star}_{jk}=\left\{\begin{array}[]{ll}{\bm{\Psi}}_{jk}&\text{if}\;\sum^{K}_{k=1}\mathrm{Tr}(\bm{\Psi}_{jk}^{H}\bm{\Psi}_{jk})\leq P\vskip 5.0pt\\ \beta_{j}\bm{\Psi}_{jk}&\text{otherwise},\end{array}\right$. (52)
@@
We only sketch the basic ideas which follow the existing work [29] closely. First, Algorithm 2 can be recognized as a MM method for problem (48); because each step of Algorithm 2 boils down to solving a convex subproblem, the convergence to a stationary point can be verified by the MM theory [29]. Moreover, using the inequality in Lemma 1 to bound $\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})$ can be thought of as constructing a surrogate function for problem (26), so each stationary point of problem (48) can be mapped to a stationary point of problem (26) according to the MM theory. ∎
Remark 5
Furthermore, our algorithms can be extended to the energy efficiency maximization. Specifically, we first decouple the ratio of the energy efficiency by Dinkelbach’s transform [7]. With an auxiliary variable iteratively updated, the new problem is the same as the problem in (26), so the proposed Algorithm 1 and Algorithm 2 can be readily applied.
Figure 1: Approximation error between the lower bounds and original objective for different noise power levels and different values of the temporal correlation coefficient $\rho_{jk}$.
@@ VI-A Settings
where $\bar{\mathbf{H}}_{jk,\ell}$ is the static component, $\mathbf{W}_{jk,\ell}\odot\mathbf{X}_{jk,\ell}$ is the random component, and $\rho_{jk}\in(0,1]$ is the tradeoff factor. These parameters are further specified in the following. The real matrix $\mathbf{W}_{jk,\ell}\in\mathbb{R}^{M_{r}\times M_{t}}$ reflects on the large-scale fading, each entry of which is computed as $10^{-\xi/10}$ where $\xi=0.5\times(128.1+37.6\log_{10}d+\tau)$, $d$ denoting the distance in km between the corresponding pair of BS and user, $\tau$ being a zero-mean Gaussian random variable with a standard deviation of 8 dB for the shadowing effect. The fixed matrix $\bar{\mathbf{H}}_{jk,\ell}$ is a realization of $\mathbf{W}_{jk,\ell}\odot\mathbf{X}_{jk,\ell}$ at the beginning of the test. The random complex matrix $\mathbf{X}_{jk,\ell}\in\mathbb{C}^{M_{r}\times M_{t}}$ has each entry independently drawn from the standard complex Gaussian distribution $\mathcal{CN}(0,1)$. The tradeoff factor $\rho_{jk}$, a.k.a. the temporal correlation coefficient, is generated by Jake’s model [13]. Specifically, we model it as $\rho_{jk}=J_{0}(2\pi v_{jk}f_{c}T/c)$, where $J_{0}(\cdot)$ is the zeroth order Bessel function of the first kind, $v_{jk}$, $f_{c}$, $c$, and $T$ are the $(j,k)$-th user’s moving speed, the carrier frequency, the speed of light, and the delay after obtaining the static component, respectively. The parameter choice in our simulations follows 3GPP_3D_UMa [1] standard: $f_{c}=4.8$ GHz, $T=0.5$ ms, $v_{jk}=60$ or $90$ km/h. The corresponding values of $\rho_{jk}$ are 0.8321 or 0.6425 respectively. We consider $1000$ independent blocks and evaluate their average rates.
@@
We further specify the network setting. The radius of each cell equals $300$ meters. We let $P=30$ dBm and $\omega_{jk}=1$. The noise power is 90-90 dBm by default. For the multi-cell case, we consider 7 wrapped-around cells. We let $M_{t}=64$, $M_{r}=2$, and $K=8$ by default for the single-cell case, and $M_{t}=32$, $K=16$, and $M_{r}=2$ by default for the multi-cell case. The proposed lower bound in Theorem 3 is compared with the existing lower bound in [4]. Moreover, the benchmark methods for stochastic precoding include:
(a) Single-cell, Rayleigh fading case
(b) Multi-cell, Rayleigh fading case
(c) Single-cell, Nakagami-$m$ fading case
(d) Multi-cell, Nakagami-$m$ fading case
Figure 2: The long-term average sum rates achieved by the different algorithms under the different fading models and under the different noise power levels when $v=60$ and $90$ km/h.
Figure 3: Convergence of different algorithms versus number of iterations under various configurations when $v=60$ km/h.
Figure 4: Convergence of different algorithms versus number of iterations under various configurations when $v=90$ km/h.
@@
Figure 5: Convergence of different algorithms versus elapsed time under various configurations when $v=60$ km/h.
@@
Figure 6: Convergence of different algorithms versus elapsed time under various configurations when $v=90$ km/h.
Figure 7: Average sum rates versus number of iterations under different algorithms or lower bounds in multi-cell system when $v=90$ km/h.
Figure 8: Convergence of different algorithms versus number of iterations under various configurations when $v=900$ km/h and users’ weights are not equal.
(a) Single-cell, Rayleigh fading case
(b) Single-cell, Nakagami-$m$ fading case
Figure 9: Convergence of different algorithms versus elapsed time under various configurations when $v=90$ km/h and users’ weights are not equal.
VI-B Test Results
We begin by comparing different lower bounds in Fig. 1. Consider the Rayleigh fading model. We remark that the proposed lower bound is evaluated with the optimal $\tilde{\bm{\Gamma}}^{\star}$ and $\tilde{\mathbf{Y}}^{\star}$, which amounts to the un-parameterized bound $\tilde{f}(\mathbf{V})$ in (23). Regarding the deterministic equivalent $\bar{f}$ in [11], we remark that it is not in closed form. More specifically, $\bar{f}$ depends on a set of auxiliary variables which remain to be updated iteratively. The approximation $\bar{f}$ equals to the original objective $f$ only after the iterative updates of these auxiliary variables converge. However, it can take infinitely many iterations to attain convergence. We have tried updating these auxiliary variables for 10 iterations (more than [11] did), but the resulting $\bar{f}$ is neither upper bound nor lower bound to $f$. For this reason, we only compare the proposed lower bound with the the existing one in [4]. It can be seen that the proposed bound is much tighter than the existing one in [4].
Next, we compare the expected sum rates achieved by different algorithms under the different values of noise power $\sigma^{2}$ in Fig. 2. For SWMMSE, we provide a total of $1000$ channel samples for online optimization. Because Algorithm 1 and Algorithm 2 only differ in the computational complexity, they produce the same performance curves in Fig. 2. Observe that the WMMSE algorithm that only depends on the static components of fading channels lead to the worst performance except for the multi-cell scenario with $v=60\text{ km/h}$. The proposed method attains much better performance. Its average sum rate is about 5% higher than that of WMMSE in the single-cell case, and 30% higher in the multi-cell case. Observe also that SWMMSE yields fairly nice performance in the single-cell case, which is almost equally good as the proposed method. However, SWMMSE becomes much worse in the presence of multiple cells since then it requires many more channel samples. In practice, online optimization by SWMMSE can be quite costly. It is noteworthy that WMMSE outperforms SWMMSE in the multi-cell scenario at $v=60\text{ km/h}$. This happens when the following two conditions are both satisfied: (i) the temporal correlation coefficient is large, so that the channels become less random and then the WMMSE is well suited for the case; (ii) there are multiple cells so that the training of the SWMMSE becomes more difficult, and thus the performance of the SWMMSE becomes worse. Thus, the WMMSE rises as the SWMMSE falls. In contrast, for the single-cell scenario, even though the performance of WMMSE still becomes better with a higher temporal correlation coefficient, the stochastic WMMSE can now be well trained and thus can yield good performance. As a result, WMMSE cannot beat the SWMMSE for the single-cell scenario even when the temporal correlation coefficient is large.